# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v2

@@ 5 Discussion and future directions
Appendix A Proof of Propositions
A.1 Proof of Proposition 1
@@ A.2 Proof of Proposition 2
where $\mathrm{(a)}$ is because $\frac{\log x}{x}$ decreases on $[e,\infty]$ and $m\geq 3000$; $\mathrm{(b)}$ is because $\log\left(\frac{1}{p^{2}\gamma}\right)=\log\left(\frac{1}{\rho p(1-p)}\right)\geq\log 4$ and $\log\left(\log\left(\frac{1}{p^{2}\gamma}\right)\right)\leq\log\left(\frac{1}{p^{2}\gamma}\right)$. Then (31) holds since $m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq m\exp\left(-\frac{mp^{2}\gamma}{192}\right)\leq m\exp(-2\log m)=m^{-1}$. We conclude the proof by applying (29) with (30) and (31).
A.3 Proof of Proposition 3
@@
where $\mathrm{(a)}$ is because $\tau_{k}\geq\frac{1}{2}\rho N_{k}$; $\mathrm{(b)}$ follows from $|\mathcal{E}_{\pi}|=N_{k}\geq\frac{mk}{3}$ and $m\rho^{2}\geq 1100\log n$. By summing (32) and (A.3) over $1\leq k\leq m$, we obtain that
@@ A.4 Proof of Proposition 4
where $\mathrm{(a)}$ is because $|\mathcal{E}_{\pi}|=N_{k}$ and $-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)={\frac{(\rho-3)|\mathcal{E}_{\pi}|}{12}}\log\left(\frac{1}{1-\rho}\right)\leq-\frac{|\mathcal{E}_{\pi}|}{6}\log\left(\frac{1}{1-\rho}\right)$; $\mathrm{(b)}$ follows from $|\mathcal{E}_{\pi}|=N_{k}=\frac{mk}{2}\left(1-\frac{k+1}{2m}\right)\geq\frac{mk}{3}$; $\mathrm{(c)}$ is because $m\geq 100\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right)$ implies $3k\log n-\frac{mk}{24}\log\frac{1}{1-\rho}\leq-k\log n$ and $\frac{k}{8}\log\frac{1}{1-\rho}-\frac{mk}{72}\log\frac{1}{1-\rho}\leq 0$. By summing over (34) and (35), we obtain that
@@
Appendix B Proof of Lemmas
B.1 Proof of Lemma 1
@@
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$ for $i=0,\dots,\ell-1$.
@@
We have two bounds for large $\ell$ in the Erdős-Rényi model, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)\leq\frac{1}{\ell}\log(\lambda_{1}^{\ell}+\lambda_{2}^{\ell})$, where $10<\alpha_{2}<\alpha_{1}<1,\alpha_{1}+\alpha_{2}=$ 1 and $\lambda_{1}>\lambda_{2}>0$. Consequently, $\lambda_{1}-\frac{\log 2}{\ell}\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq\lambda_{1}+\frac{\log 2}{\ell}.$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $\|x\|_{\ell}\leq\|x\|_{2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
@@
where $(\mathbf{W}_{\ell})_{00}=1,(\mathbf{W}_{\ell})_{ii}=1-t^{2}(1-\rho^{2})$ for any $1\leq i\leq\ell$ and $(\mathbf{W}_{\ell})_{ij}=-t\rho$ for any $|i-j|=1$ with $0\leq i,j\leq\ell$. By (43), we have
@@
Since $01-t^{2}(1-\rho^{2})-2t\rho=(1-t\rho-t)(1-t\rho+t)>$ 0 by $t<\frac{1}{1+\rho}$, we obtain $(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}>0$, and hence
@@
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for $i=0,\dots,\ell-1$. Recall that $\lambda_{1}+\lambda_{2}=1-t^{2}(1-\rho^{2})$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}$. Following a similar argument as (43), we have
@@
where $\mathrm{(a)}$ is because $t^{2}\rho^{2}<\frac{\rho^{2}}{(1+\rho)^{2}}<1$ and $12+2t^{2}>$ 1.
@@
where $(\mathbf{W}_{\ell})_{00}=1+t,(\mathbf{W}_{\ell})_{\ell\ell}=1+t(2-\rho^{2}),(\mathbf{W}_{\ell})_{ii}=1+2t$ for any $11\leq i\leq\ell-$ 1 and $(\mathbf{W}_{\ell})_{ij}=-t\rho$ for any $|i-j|=1$ with $0\leq i,j\leq\ell$. We note that the symmetric matrix $\mathbf{W}_{\ell}$ is strictly diagonally dominant with positive diagonal entries, and then it is positive definite. This follows from the eigenvalues of symmetric matrix being real, and the Gershgorin’s circle theorem [Ger31]. Denote $\mathbf{W}_{\ell}=\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}$ with $\mathbf{J}_{\ell}\in\mathbb{R}^{(\ell+1)\times(\ell+1)}$. Let $C_{i}=\frac{B_{i}}{\sqrt{1+t(1-\rho^{2})}}$ for all $0\leq i\leq\ell$ and denote $\mathbf{C}_{\ell}\triangleq[C_{0},C_{1},\cdots,C_{\ell}]$. By (47), we note that
@@
where $(\mathbf{U}_{\ell})_{00}=1+t,(\mathbf{U}_{\ell})_{ii}=1+2t$ for any $1\leq i\leq\ell$ and $(\mathbf{U}_{\ell})_{ij}=-t\rho$ for any $|i-j|=1$ with $0\leq i,j\leq\ell$. Expanding the last column of $\mathbf{W}_{\ell}$ and $\mathbf{U}_{\ell}$ yields that
@@
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for $i=0,\dots,\ell-1$. Recall that $\lambda_{1}+\lambda_{2}=1+2t$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}$. Following a similar argument as (47), we have
@@
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell-1}(1+2t)C_{i}^{2}+\frac{1}{2}\sum_{i=0}^{\ell-1}t\rho C_{i}C_{i+1}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell-1}$
@@
We have two bounds for large $\ell$ in the Gaussian Wigner model, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)\geq\frac{1}{\ell}\log\left(\left(\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\right)^{2}\right)$, where $\alpha_{2}<0<1<\alpha_{1},\alpha_{1}+\alpha_{2}=1$ and $\lambda_{1}>\lambda_{2}>0$. Consequently, $-2\left(\lambda_{1}+\frac{\log\alpha_{1}}{\ell}\right)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq-2\left(\lambda_{1}+\frac{1}{\ell}\log\left[1-2\left(\frac{\lambda_{2}}{\lambda_{1}}\right)^{\ell/2}\right]\right).$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $(x+y)^{\ell/2}-y^{\ell/2}\geq x^{\ell/2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
B.2 Proof of Lemma 2

### v3

@@ 5 Discussion and future directions
Appendix A Support Recovery Problem
We prove the impossibility results for exact recovery of $(S^{*},T^{*})$ by Fano’s method. On the one hand, since $\pi^{*}$ contains the information of $S^{*}$ and $T^{*}$, the mutual information can be upper bounded by
$I(S^{*},T^{*};G_{1},G_{2})\leq I(\pi^{*};G_{1},G_{2})\leq 25\binom{m}{2}p^{2}\phi(\gamma).$
On the other hand, for exact recovery, $\mathcal{M}$ is simply the set of all $(S,T)$, and thus $|\mathcal{M}|=\binom{n}{m}^{2}$. Applying the Fano’s inequality, we obtain that
$\mathbb{P}\left[(\hat{S},\hat{T})\neq(S^{*},T^{*})\right]\geq 1-\frac{I(S^{*},T^{*};G_{1},G_{2})+\log 2}{\log|\mathcal{M}|}\geq 1-\frac{25\binom{m}{2}p^{2}\phi(\gamma)+\log 2}{2\log\binom{n}{m}}$.
Since $\binom{n}{m}\geq\left(\frac{n}{m}\right)^{m}$, the impossibility results follow if
$25mp^{2}\phi(\gamma)\leq 4c(\log n-\log m).$ (25)
Consider the regime that $\frac{\log n}{p^{2}\phi(\gamma)}\geq 13\left(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\log\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\right)$. We first show $\frac{\log n}{p^{2}\phi(\gamma)}=O(n^{1-\delta_{0}})$ for some constant $\delta_{0}>0$. Indeed, if $\frac{\log n}{p^{2}\phi(\gamma)}=n^{1-o(1)}$, we have $p^{2}\phi(\gamma)=n^{-1+o(1)}$, and thus $2\log(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})})\geq\log n$. When $\gamma\geq 1$, we have $\phi(\gamma)\geq\frac{\gamma}{6}$. When $0<\gamma<1$, we have $\phi(\gamma)\geq\frac{\gamma^{2}}{6}$. Therefore, $p^{2}\phi(\gamma)\geq\frac{1}{6}p^{2}(\gamma\wedge\gamma^{2})$. We obtain that $\frac{\log n}{p^{2}\phi(\gamma)}\leq\frac{12\log(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})})}{p^{2}(\gamma\wedge\gamma^{2})}$, which is contradictory with $\frac{\log n}{p^{2}\phi(\gamma)}\geq 13\left(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\log\left(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\right)\right)$. Consequently, $\frac{\log n}{p^{2}\phi(\gamma)}=O(n^{1-\delta_{0}})$ for some $\delta_{0}>0$. We note that $m\leq\frac{4c\delta_{0}\log n}{25p^{2}\phi(\gamma)}$ suffices for (25), and thus proves the impossibility results.
Appendix B Proof of Propositions
B.1 Proof of Proposition 1
@@ B.2 Proof of Proposition 2
where $\mathrm{(a)}$ is because $\frac{\log x}{x}$ decreases on $e,\infty]$ and $m\geq 3000$; $\mathrm{(b)}$ is because $\log\left(\frac{1}{p^{2}\gamma}\right)=\log\left(\frac{1}{\rho p(1-p)}\right)\geq\log 4$ and $\log\left(\log\left(\frac{1}{p^{2}\gamma}\right)\right)\leq\log\left(\frac{1}{p^{2}\gamma}\right)$. Then (33) holds since $m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq m\exp\left(-\frac{mp^{2}\gamma}{192}\right)\leq m\exp(-2\log m)=m^{-1}$. We conclude the proof by applying (31) with (32) and (33).
B.3 Proof of Proposition 3
@@
where $\mathrm{(a)}$ is because $\tau_{k}\geq\frac{1}{2}\rho N_{k}$; $\mathrm{(b)}$ follows from $|\mathcal{E}_{\pi}|=N_{k}\geq\frac{mk}{3}$ and $m\rho^{2}\geq 1100\log n$. By summing (34) and (B.3) over $1\leq k\leq m$, we obtain that
@@ B.4 Proof of Proposition 4
where $\mathrm{(a)}$ is because $|\mathcal{E}_{\pi}|=N_{k}$ and $\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)={\frac{(\rho-3)|\mathcal{E}_{\pi}|}{12}}\log\left(\frac{1}{1-\rho}\right)\leq-\frac{|\mathcal{E}_{\pi}|}{6}\log\left(\frac{1}{1-\rho}\right)$; $\mathrm{(b)}$ follows from $|\mathcal{E}_{\pi}|=N_{k}=\frac{mk}{2}\left(1-\frac{k+1}{2m}\right)\geq\frac{mk}{3}$; $\mathrm{(c)}$ is because $m\geq 100\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right)$ implies $3k\log n-\frac{mk}{24}\log\frac{1}{1-\rho}\leq-k\log n$ and $\frac{k}{8}\log\frac{1}{1-\rho}-\frac{mk}{72}\log\frac{1}{1-\rho}\leq 0$. By summing over (36) and (37), we obtain that
@@
Appendix C Proof of Lemmas
C.1 Proof of Lemma 1
@@
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $e_{1}\dots e_{\ell}]$ as illustrated in Figure 5. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$ for $i=0,\dots,\ell-1$.
@@
We have two bounds for large $\ell$ in the Erdős-Rényi model, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)\leq\frac{1}{\ell}\log(\lambda_{1}^{\ell}+\lambda_{2}^{\ell})$, where $0<\alpha_{2}<\alpha_{1}<1,\alpha_{1}+\alpha_{2}=1$ and $\lambda_{1}>\lambda_{2}>0$. Consequently, $\lambda_{1}-\frac{\log 2}{\ell}\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq\lambda_{1}+\frac{\log 2}{\ell}.$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $\|x\|_{\ell}\leq\|x\|_{2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
@@
where $(\mathbf{W}_{\ell})_{00}=1,(\mathbf{W}_{\ell})_{ii}=1-t^{2}(1-\rho^{2})$ for any $1\leq i\leq\ell$ and $(\mathbf{W}_{\ell})_{ij}=-t\rho$ for any |ij|=1|i-j|=1 with $0\leq i,j\leq\ell$. By (45), we have
@@
Since $1-t^{2}(1-\rho^{2})-2t\rho=(1-t\rho-t)(1-t\rho+t)>0$ by $t<\frac{1}{1+\rho}$, we obtain $(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}>0$, and hence
@@
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $e_{1}\dots e_{\ell}]$ as illustrated in Figure 5. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for $i=0,\dots,\ell-1$. Recall that $\lambda_{1}+\lambda_{2}=1-t^{2}(1-\rho^{2})$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}$. Following a similar argument as (45), we have
@@
where $\mathrm{(a)}$ is because $t^{2}\rho^{2}<\frac{\rho^{2}}{(1+\rho)^{2}}<1$ and $2+2t^{2}>1$.
@@
where $(\mathbf{W}_{\ell})_{00}=1+t,(\mathbf{W}_{\ell})_{\ell\ell}=1+t(2-\rho^{2}),(\mathbf{W}_{\ell})_{ii}=1+2t$ for any $1\leq i\leq\ell-1$ and $(\mathbf{W}_{\ell})_{ij}=-t\rho$ for any |ij|=1|i-j|=1 with $0\leq i,j\leq\ell$. We note that the symmetric matrix $\mathbf{W}_{\ell}$ is strictly diagonally dominant with positive diagonal entries, and then it is positive definite. This follows from the eigenvalues of symmetric matrix being real, and the Gershgorin’s circle theorem [Ger31]. Denote $\mathbf{W}_{\ell}=\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}$ with $\mathbf{J}_{\ell}\in\mathbb{R}^{(\ell+1)\times(\ell+1)}$. Let $C_{i}=\frac{B_{i}}{\sqrt{1+t(1-\rho^{2})}}$ for all $0\leq i\leq\ell$ and denote $\mathbf{C}_{\ell}\triangleq[C_{0},C_{1},\cdots,C_{\ell}]$. By (49), we note that
@@
where $(\mathbf{U}_{\ell})_{00}=1+t,(\mathbf{U}_{\ell})_{ii}=1+2t$ for any $1\leq i\leq\ell$ and $(\mathbf{U}_{\ell})_{ij}=-t\rho$ for any |ij|=1|i-j|=1 with $0\leq i,j\leq\ell$. Expanding the last column of $\mathbf{W}_{\ell}$ and $\mathbf{U}_{\ell}$ yields that
@@
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $e_{1}\dots e_{\ell}]$ as illustrated in Figure 5. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for $i=0,\dots,\ell-1$. Recall that $\lambda_{1}+\lambda_{2}=1+2t$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}$. Following a similar argument as (49), we have
@@
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell-1}(1+2t)C_{i}^{2}+\sum_{i=0}^{\ell-1}t\rho C_{i}C_{i+1}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell-1}$
@@
We have two bounds for large $\ell$ in the Gaussian Wigner model, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)\geq\frac{1}{\ell}\log\left(\left(\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\right)^{2}\right)$, where $\alpha_{2}<0<1<\alpha_{1},\alpha_{1}+\alpha_{2}=1$ and $\lambda_{1}>\lambda_{2}>0$. Consequently, $2\left(\lambda_{1}+\frac{\log\alpha_{1}}{\ell}\right)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq-2\left(\lambda_{1}+\frac{1}{\ell}\log\left[1-2\left(\frac{\lambda_{2}}{\lambda_{1}}\right)^{\ell/2}\right]\right).$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $(x+y)^{\ell/2}-y^{\ell/2}\geq x^{\ell/2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
C.2 Proof of Lemma 2